# The Assouad dimension of Kakeya sets in $\mathbb{R}^{3}$

### v1

@@ 4.2 Arrangements of tubes that cluster into rectangular prisms
We would like to apply Lemma 4.8 to each set $\mathbb{T}^{\dagger}_{s_{N}},\ldots,\mathbb{T}^{\dagger}_{s_{1}}$ in turn. If Conclusion (B) of Lemma 4.8 holds at any step, then we conclude that Conclusion (B) of Lemma 4.10 holds and we halt. If not, then few tubes from $\mathbb{T}^{\dagger}_{s_{i}}$ can cluster into a $s_{i}\times t\times 1$ prism; if each tube from $\mathbb{T}^{\dagger}_{s_{i}}$ contains the same number of tubes from $\mathbb{T}$, then Conclusion (A) of Lemma 4.10 holds, and we are done. The difficulty is that conclusion (A) of Lemma 4.10 only holds for a subset of the tubes in $\mathbb{T}^{\dagger}_{s_{i}}$; this complication introduces additional pigeonholing at each step, which in turn forces us to re-label all of the relevant objects. We now turn to the unpleasant details.
Define $\mathbb{T}^{(1)}_{s_{N}}=\mathbb{T}^{{\dagger}}_{s_{N}}$ and apply Lemma 4.8 with $\varepsilon/2$ in place of $\varepsilon$ to $\mathbb{T}^{(1)}_{s_{N}}$, with the shading $Y$ inherited from $\mathbb{T}$ (recall that this shading is uniformly $\delta^{\eta}$ dense). If Conclusion (B) of Lemma 4.8 holds, then Conclusion (B) of Lemma 4.10 holds, and we are done.
If Conclusion (A) of Lemma 4.8 holds, then we obtain a set $(\mathbb{T}^{(1)}_{s_{N}})^{\prime}\subset\mathbb{T}^{(1)}_{s_{N}}$ with #$\#(\mathbb{T}^{(1)}_{s_{N}})^{\prime}\geq\frac{1}{2}(\#\mathbb{T}^{(1)}_{s_{N}})$, so that for each $t\in[\delta,1]$ and each $2s_{N}\times 2t\times 2$ rectangle $R$ (recall $s_{N}=\delta$), at most $\delta^{-\varepsilon/2}(t/\delta)$ tubes from $(\mathbb{T}^{(1)}_{s_{N}})^{\prime}$ are contained in $R$. After dyadic pigeonholing and refining $(\mathbb{T}^{(1)}_{s_{N}})^{\prime}$, we can select sets $\mathbb{T}^{(2)}_{s_{N}}\subset(\mathbb{T}^{(1)}_{s_{N}})^{\prime}$ and $\mathbb{T}^{(2)}_{s_{N-1}}\subset\mathbb{T}^{{\dagger}}_{s_{N-1}}$ so that #$\#\mathbb{T}^{(2)}_{s_{N}}\gtrsim|\log\delta|^{-1}(\mathbb{T}^{(1)}_{s_{N}})^{\prime}$, and #$\#(\mathbb{T}^{(2)}_{s_{N}}[T_{s_{N-1}}]$ is the same for each $T_{s_{N-1}}\in\mathbb{T}^{(2)}_{s_{N-1}}$, i.e. $\mathbb{T}^{(2)}_{s_{N-1}}$ is a 1-balanced partitioning cover of $\mathbb{T}^{(2)}_{s_{N}}$.
Next, we will describe a slightly more elaborate version of the above process. Let $2\leq k\leq N$, and suppose we have constructed sets $\mathbb{T}_{s_{N}}^{(k)},\ldots,\mathbb{T}_{s_{N-k+1}}^{(k)}$, that have the following properties:
  • (a)
    For $N-k+1\leq i<N$, $\mathbb{T}_{s_{i}}^{(k)}$ is a 1-balanced partitioning cover of $\mathbb{T}_{s_{i+1}}^{(k)}$ (and hence also a 1-balanced partitioning cover of $\mathbb{T}_{s_{j}}^{(k)}$ for $j>i$).
  • (b)
    #$\#\mathbb{T}_{s_{N}}^{(k)}\gtrsim_{k}|\log\delta|^{-k}(\#\mathbb{T}_{s_{N}}^{(1)})$.
  • (c)
    For $N-k+2\leq i\leq N$ and $s_{i}\leq t\leq 1$, each $s_{i}\times t\times 1$ rectangle $R$ contains $\lesssim_{k}|\log\delta|^{k}\delta^{-\varepsilon/2}\frac{\#\mathbb{T}_{s_{N}}^{(k)}}{\#\mathbb{T}_{s_{i}}^{(k)}}$ tubes from $\mathbb{T}_{N}^{(k)}$.
We construct sets $\mathbb{T}_{s_{N}}^{(k+1)},\ldots,\mathbb{T}_{s_{N-k}}^{(k+1)}$ as follows. For each $T_{s_{N-k+1}}\in\mathbb{T}^{(k)}_{s_{N-k+1}}$, define the shading
$Y(T_{N-k+1})=T_{s_{N-k+1}}\cap N_{s_{N-k+1}}\Big{(}\bigcup_{T\in\mathbb{T}}Y(T)\Big{)}.$
Since $\mathbb{T}[T_{s_{N-k+1}}]$ is non-empty and each tube $T\in\mathbb{T}$ has a $\delta^{\eta}$-dense shading, we conclude that $|Y(T_{s_{N-k+1}})|\gtrsim\delta^{\eta}|T_{s_{N-k+1}}|$, i.e. the shading is uniformly $\gtrsim\delta^{\eta}$-dense. Since $s_{N-k+1}\geq\delta^{1/N}$, this means that the shading is uniformly $s_{N-k+1}^{\eta N}$ dense. We will select $\eta$ and $\delta_{0}$ sufficiently small so that Lemma 4.8 can be applied to the pair $(\mathbb{T}^{(k)}_{s_{N-k+1}},Y)_{s_{N-k+1}}$ with $s_{N-k+1}$ in place of $\delta$, and $\varepsilon/2$ in place of $\varepsilon$. Let $\eta_{1}=\eta_{1}(\varepsilon)>0$ be the output of that lemma.
@@
If Conclusion (B) of Lemma 4.10 holds, then $\bigcup_{T_{s_{N-k+1}}\in\mathbb{T}^{(k)}_{s_{N-k+1}}}Y(T_{s_{N-k+1}})$ has discretized Assouad dimension at least $23-\varepsilon/$ 2 at scale $s_{N-k+1}$ and scale separation $s_{N-k+1}^{-\eta_{1}}$. But this means that $\bigcup_{T\in\mathbb{T}}Y(T)$ has has discretized Assouad dimension at least $23-\varepsilon/$ 2 at scale $\delta$ and scale separation $\delta^{-\eta_{1}/N}$; if we select $\eta\leq\eta_{1}/N$, then Conclusion (B) of Lemma 4.10 holds, and we are done.
Suppose instead that Conclusion (A) of Lemma 4.10 holds. Then there exists a set $(\mathbb{T}^{(k)}_{s_{N-k+1}})^{\prime}\subset\mathbb{T}^{(k)}_{s_{N-k+1}}$ with #$\#(\mathbb{T}^{(k)}_{s_{N-k+1}})^{\prime}\geq\frac{1}{2}(\#\mathbb{T}^{(k)}_{s_{N-k+1}})$, so that for each $s_{N-k+1}\leq t\leq 1$ and each $2s_{N-k+1}\times 2t\times 2$ rectangle $R$, at most $s_{N-k+1}^{-\varepsilon/2}\frac{t}{s_{N-k+1}}\leq\delta^{-\varepsilon/2}\frac{t}{s_{N-k+1}}$ tubes from $(\mathbb{T}^{(k)}_{s_{N-k+1}})^{\prime}$ are contained in $R$.
After dyadic pigeonholing and refining $(\mathbb{T}^{(k)}_{s_{N-k+1}})^{\prime}$, we can select sets $\mathbb{T}^{(k+1)}_{s_{N-k+1}}\subset(\mathbb{T}^{(k)}_{s_{N-k+1}})^{\prime}$ and $\mathbb{T}^{(k+1)}_{s_{N-k}}\subset\mathbb{T}^{{\dagger}}_{s_{N-k}}$ so that #$\#\mathbb{T}^{(k+1)}_{s_{N-k+1}}\gtrsim|\log\delta|^{-1}(\mathbb{T}^{(k)}_{s_{N-k+1}})^{\prime}$, and $\mathbb{T}^{(k+1)}_{s_{N-k}}$ is a 1-balanced partitioning cover of $\mathbb{T}^{(k+1)}_{s_{N-k+1}}$. For each $N-k+1<i\leq N$, define $\mathbb{T}^{(k+1)}_{s_{i}}$ to be those tubes from $\mathbb{T}^{(k)}_{s_{i}}$ that are contained in some tube from $\mathbb{T}^{(k+1)}_{s_{N-k+1}}$; it is still the case that $\mathbb{T}_{s_{i}}$ is a 1-balanced partitioning cover of $\mathbb{T}_{s_{i+1}}$, for $N-k+1\leq i<N$, and by construction it is also the case that $\mathbb{T}^{(k+1)}_{s_{N-k}}$ is a 1-balanced partitioning cover of $\mathbb{T}^{(k+1)}_{s_{N-k+1}}$.
Furthermore, #$\#\mathbb{T}^{(k+1)}_{s_{N}}\gtrsim|\log\delta|^{-1}(\#\mathbb{T}^{(k)}_{s_{N}})$ (this is because the cardinality of $\mathbb{T}^{(k)}_{s_{N-k+1}}$ was reduced by a factor of $\lesssim|\log\delta|^{-1}$, and $\mathbb{T}^{(k)}_{s_{N-k+1}}$ was a 1-balanced partitioning cover of $\mathbb{T}^{(k)}_{s_{N}}$), and thus Item (c) remains true for $k+1$ in place of $j$, and Item (c) remains true for $N-k+2\leq i\leq N$, with $k+1$ in place of $k$ (indeed, the cardinality of $\mathbb{T}_{s_{N}}^{(k+1)}$ has decreased by at most a multiplicative factor of $O(|\log\delta|)$ compared to the cardinality of $\mathbb{T}_{s_{N}}^{(k)}$, and the cardinality of $\mathbb{T}_{s_{i}i}^{(k+1)}$ is no larger than that of $\mathbb{T}_{s_{i}i}^{(k)}$)
All that remains is to verify Item (c) above for $i=N-k+1$. Let $s_{N-k+1}\leq t\leq 1$, let $R$ be a $s_{N-k+1}\times t\times 1$ rectangle, and let $R^{\prime}=2R$. We already know that at most $\delta^{-\varepsilon/2}\frac{t}{s_{N-k+1}}$ tubes from $\mathbb{T}^{(k+1)}_{s_{N-k+1}}$ can be contained in the $R^{\prime}$. On the other hand, if $T$ is a $\delta$-tube that is contained in $R$ and also contained in a $s_{n-k+1}$ tube $T_{s_{N-k+1}}$, then $T_{s_{N-k+1}}\subset 2R$. Since $\mathbb{T}^{(k+1)}_{s_{N-k+1}}$ is a 1-balanced cover of $\mathbb{T}^{(k+1)}_{s_{N}}$, each tube in $\mathbb{T}^{(k+1)}_{s_{N-k+1}}$ contains #$\frac{\#\mathbb{T}^{(k+1)}_{s_{N}}}{\#\mathbb{T}^{(k+1)}_{s_{N-k+1}}}$ tubes from $\mathbb{T}^{(k+1)}_{s_{N}}$. We conclude that $\lesssim\delta^{-\varepsilon/2}\frac{t}{s_{N-k+1}}\frac{\#\mathbb{T}^{(k+1)}_{s_{N}}}{\#\mathbb{T}^{(k+1)}_{s_{N-k+1}}}$ tubes from $\mathbb{T}^{(k+1)}_{s_{N}}$ are contained in $R$. The completes the $k\to k+1$ step of the above iteration.
@@
If the iterative process described above does not halt before the $N$-th, step, then define $\mathbb{T}^{\prime}=\mathbb{T}^{(N)}_{s_{N}}$. By selecting $\delta_{0}$ sufficiently small we ensure that #$\#\mathbb{T}^{\prime}\geq\delta^{-\varepsilon}(\#\mathbb{T})$. For each $s=\delta^{i/N},\ i=1,\ldots,N$, we have the 1-balanced partitioning cover $\mathbb{T}^{\prime}_{s}=\mathbb{T}^{(N)}_{s_{i}}$. If we select $\delta_{0}$ sufficiently small depending on $\varepsilon$, then Item (c) above ensures that for every $t\in[s,1]$ and every $s\times t\times 1$ rectangular prism $R$, at most $\delta^{-\varepsilon}\frac{t}{s}\frac{(\#\mathbb{T}^{\prime})}{(\#\mathbb{T}^{\prime}_{s})}$ tubes from $\mathbb{T}^{\prime}$ are contained in $R$. ∎
@@
For all $\varepsilon>0$, there exists $\tau>0$, so that for all $\eta_{1}>0$, there exists $\eta,\delta_{0}>0$ so that the following holds for all $\delta\in(0,\delta_{0}]$. Let $\beta\geq 2$ and let $(\mathbb{T},Y)_{\delta}$ be a set of $\delta^{-\beta}$ essentially distinct $\delta$-tubes that satisfy the Convex Wolff Axioms with error $\delta^{\eta}$, with $Y$ a $\delta^{\eta}$ uniformly dense shading. Then at least one of the following must be true.
@@
    $\bigcup Y(T)$ has discretized Assouad dimension at least $3-\varepsilon$ at scale $\delta$, with scale separation $\delta^{-\eta}$.
@@
    There exists $\rho\in[\delta,\delta^{\varepsilon}]$ and a set $\tilde{\mathbb{T}}$ of $\rho$-tubes, with the following properties.
@@
    There exists $\rho\in[\delta^{1-\varepsilon},1]$, a $\rho$-tube $T_{\rho}$, and a set $\tilde{\mathbb{T}}\subset\mathbb{T}[T_{\rho}]$, so that the following holds.
@@
Let $\eta_{2}=1/N>0$ be a small quantity to be chosen below. Apply Lemma 4.10 with $\eta_{2}$ in place of $\varepsilon$. We conclude that either Conclusion (A) of Proposition 4.1 holds, there exists a set $\mathbb{T}^{\prime}\subset\mathbb{T}$ that satisfies Conclusion (B) of Lemma 4.10.
Step 2.
Suppose there exists $\rho\in[\delta,\delta^{\varepsilon}]$ of the form $\rho=\delta^{i/N}$ so that #$\#\mathbb{T}^{\prime}_{\rho}\geq\delta^{-\beta-\tau}$ (here #$\#\mathbb{T}^{\prime}_{\rho}$ is a 1-uniform partitioning cover of $\mathbb{T}^{\prime}$, coming from Conclusion (B) of Lemma 4.10). Then by Remark 2.11, $\mathbb{T}^{\prime}_{\rho}$ satisfies the Convex Wolff Axioms with error $\delta^{-2\eta}\leq\rho^{-2\eta/\varepsilon}$; this is at most $\rho^{-\eta_{1}}$, provided we choose $\eta\leq\varepsilon\eta_{1}/2$. Thus Conclusion (C) holds, with $\tilde{\mathbb{T}}=\mathbb{T}^{\prime}_{\rho}$.
Step 3. Our goal in this step is to show that either Conclusion (D) of Proposition 4.1 holds, or for every $\rho\in[\delta,\delta^{\varepsilon}]$ of the form $\rho=\delta^{i/N}$, we have
$\rho^{-\beta+\varepsilon/3}\leq\#\mathbb{T}^{\prime}_{\rho}\leq\rho^{-\beta-\tau},\quad\rho\in[\delta,\delta^{\varepsilon}].$ (4.31)
Note that the RHS of the above inequality was already established in Step 2. First, we consider the case where $\beta\leq 2+\varepsilon/4$; in this case (4.31) will always hold. Indeed, let $\rho\in[\delta,\delta^{\varepsilon}]$ be of the form $\rho=\delta^{i/N}$. Then for each $T_{\rho}\in\mathbb{T}_{\rho}^{\prime}$, we have
#$\frac{\#\mathbb{T}^{\prime}}{\#\mathbb{T}^{\prime}_{\rho}}=\#\mathbb{T}^{\prime}[T_{\rho}]\leq\delta^{-2\eta}\rho^{2}\#\mathbb{T}^{\prime},$
where the second inequality follows from the fact that $\mathbb{T}^{\prime}$ satisfies the Convex Wolff Axioms with error $\delta^{-2\eta}$. We conclude that #$\#\mathbb{T}^{\prime}_{\rho}\geq\delta^{2\eta}\rho^{-2}$. In particular, if $\beta\leq 2+\varepsilon/4$, and if we select $\eta$ sufficiently small, then
#$\#\mathbb{T}^{\prime}_{\rho}\geq\delta^{2\eta}\rho^{-2}\geq\rho^{2\eta/\varepsilon+\varepsilon/4}\rho^{-\beta}\geq\rho^{-\beta+\varepsilon/3},$ (4.32)
@@
Next we consider the case where $\beta>2+\varepsilon/4$. Suppose there exists some $\rho\in[\delta,\delta^{\varepsilon}]$ of the form $\rho=\delta^{i/N}$ for which #$\#\mathbb{T}^{\prime}_{\rho}<\rho^{-\beta+\varepsilon/4}.$ Let $\rho\in[\delta,\delta^{\varepsilon}]$ be the smallest number of the form $\delta^{iN}$ for which this is true. Note that $\rho\geq\delta^{1-\varepsilon/20}$, since the tubes in $\mathbb{T}^{\prime}$ are essentially distinct, and thus #$\#\mathbb{T}^{\prime}_{\rho}\geq(\delta/\rho)^{4}(\#\mathbb{T}^{\prime})\geq(\delta/\rho)^{4}\delta^{-\beta+2\eta}$.
Let $T_{\rho}\in\mathbb{T}_{\rho}$. By Conclusion (B) of Lemma 4.10, for each $s\in[\delta,\rho]$ of the form $\delta^{jN}$, each $t\in[s,\rho]$ and each $s\times t\times 1$ rectangular prism $R\subset T_{\rho}$, we have
#$\#\{T\in\mathbb{T}^{\prime}[T_{\rho}]\colon T\subset R\}$ $\leq\delta^{-\eta_{2}}\frac{t}{s}\frac{(\#\mathbb{T}^{\prime})}{(\#\mathbb{T}^{\prime}_{s})}\leq\delta^{-\eta_{2}}\frac{t}{s}\frac{\delta^{-\beta}}{\rho^{-\beta+\varepsilon/4}}\Big{(}\frac{s}{\rho}\Big{)}^{\beta-\varepsilon/4}$
@@
where the second inequality used the fact that #$\#\mathbb{T}_{s}^{\prime}\geq s^{-\beta+\varepsilon/4}$ and the minimality of $\rho$, the third inequality used the assumption $\beta-\varepsilon/4\geq 2$, and the fourth inequality used the assumptions #$\#\mathbb{T}^{\prime}\geq\delta^{-\beta+\eta_{1}}$ and #$\#\mathbb{T}_{\rho}\leq\rho^{-\beta+\varepsilon/4}$.
@@
Define $\tilde{\mathbb{T}}=\mathbb{T}^{\prime}[T_{\rho}]$, and let $W\subset T_{\rho}$ be a convex set that contains at least one tube from $\tilde{\mathbb{T}}$. Then $W$ is comparable to a $s\times t\times 1$ prism, for some $\delta\leq s\leq t\leq\rho$; increasing $s$ (and possibly $t$ by a factor of at most $\delta^{-1/N}$, we may suppose that $s$ is of the form $\delta^{jN}$. Thus we have
@@
#$\#\{T\in\tilde{\mathbb{T}}\colon T\subset W\}\lesssim\delta^{-4\eta_{2}}\frac{|W|}{|T_{\rho}|}(\#\tilde{\mathbb{T}})\leq(\delta/\rho)^{-40\eta_{2}/\varepsilon}\frac{|W|}{\rho^{2}}(\#\tilde{\mathbb{T}}).$
@@
If we select $\eta_{1}<40\eta_{2}\varepsilon$, then our set $\tilde{\mathbb{T}}$ satisfies Conclusion (D) from Proposition 4.1 and we are done.
@@
Step 4. At this point, we have reduced to the case where (4.31) holds for all $\rho\in[\delta,\delta^{\varepsilon}]$ of the form $\rho=\delta^{i/N}$. We will show that Conclusion (B) holds. To begin, note that Items (i) and (iii) from Definition 2.15 hold for $\mathbb{T}^{\prime}$: for each $\rho_{0}\in[\delta,\delta^{\varepsilon}]$, we can select $\rho\in[\rho_{0},\delta^{-\varepsilon}\rho_{0}]$ of the form $\rho=\delta^{i/N}$; $\mathbb{T}_{\rho}$ is a partitioning cover of $\mathbb{T}^{\prime}$, so Item (i) holds, and Item (iii) holds by (4.31) (if $\rho_{0}>\delta^{\varepsilon}$ then we can select $\rho=1$ and there is nothing to prove).

### v2

@@ 4.2 Arrangements of tubes that cluster into rectangular prisms
For each $i=1,\ldots,N$, apply Lemma 4.7 to $(\mathbb{T}_{s_{i}},Y_{s_{i}})_{s_{i}}$ with $\varepsilon$ as above; we can do this provided $\eta>0$ and $\delta_{0}$ are selected sufficiently small depending on $\varepsilon$ (recall that $N=1/\varepsilon$). If Conclusion (A) of Lemma 4.7 holds, then $\bigcup_{\mathbb{T}_{s_{i}}}Y_{s_{i}}(T_{s_{i}})$ has discretized Assouad dimension at least $3-\varepsilon$ at scale $s_{i}$ and scale separation $s_{i}^{-\eta^{\prime}}$, for some $\eta^{\prime}$ depending only on $\varepsilon$. This in turn implies that $\bigcup_{\mathbb{T}}Y(T)$ has discretized Assouad dimension at least $3-\varepsilon$ at scale $\delta$ (this is true since $\delta\leq s_{i}$) and scale separation $\delta^{-\eta^{\prime}/N}$. Thus Conclusion (A) of Lemma 4.9 holds, provided we select $\eta\leq\eta^{\prime}/N=\eta^{\prime}\varepsilon$. If this happens then we are done.
Suppose instead that Conclusion (B) of Lemma 4.7 holds. This gives us a set $\mathbb{T}_{s_{i}}^{\prime}\subset\mathbb{T}_{s_{i}}$ with #$\#\mathbb{T}_{s_{i}}^{\prime}\geq\frac{1}{2}(\#\mathbb{T}_{s_{i}})$ so that for every $s_{i}\leq t\leq 1$ and every $s_{i}\times t\times 1$ rectangular prism $R$, at most $s_{i}^{-\varepsilon}(t/s_{i})\leq\delta^{-\varepsilon}(t/s_{i})$ tubes from $\mathbb{T}_{s_{i}}^{\prime}$ are contained in $100R$. Abusing notation slightly, we will replace each set $\mathbb{T}_{s_{j}},\ j>i$ with $\bigcup_{\mathbb{T}_{s_{i}}^{\prime}}\mathbb{T}_{s_{j}}[T_{s_{i}}]$, i.e. we only keep those tubes from $\mathbb{T}_{s_{j}}$ that are contained in some tube from $\mathbb{T}_{s_{i}}^{\prime}$. This process discards at most half the tubes from $\mathbb{T}_{s_{j}}$, so at each stage in this process, $\mathbb{T}_{s_{j}}$ still satisfies the convex Wolff axioms with error at most $2^{N}s_{j}^{-2\eta N}$ and hence we can continue to iterate Step 2.
@@
Suppose that Conclusion (B) of Lemma 4.7 holds for each $i=1,\ldots,N$ in Step 2. While each set $\mathbb{T}_{s_{i}}$ is a uniform partitioning cover of $\mathbb{T}^{\dagger}$, this property might no longer hold for the sets $\mathbb{T}_{s_{i}}^{\prime}$. We will fix this with another round of pigeonholing. This gives us sets $\mathbb{T}_{s_{i}}^{\prime}\subset\mathbb{T}_{s_{i}}$, $i=1,\ldots,N$ and a set $\mathbb{T}^{\prime}\subset\mathbb{T}^{\dagger}$, so that #$\#\mathbb{T}^{\prime}\geq(100\log(1/\delta))^{-N}(\#\mathbb{T}^{\dagger})$; each set $\mathbb{T}_{s_{i}}^{\prime}$ is a uniform partitioning cover of $\mathbb{T}^{\prime}$; and for each index $i$ and each $s_{i}\times t\times 1$ prism $R$ we have
#$\#\mathbb{T}^{\prime}[100R]\leq\#\mathbb{T}_{s_{i}}^{\prime}[100R]\frac{\#\mathbb{T}^{\prime}}{\#\mathbb{T}_{s_{i}}^{\prime}}\leq\delta^{-\varepsilon}\frac{t}{s_{i}}\frac{(\#\mathbb{T}^{\prime})}{(\#\mathbb{T}_{s_{i}}^{\prime})}.$
Thus $\mathbb{T}^{\prime}$ and $\mathbb{T}_{s_{i}}^{\prime},\ i=1,\ldots,N$ satisfy Conclusion (B) of Lemma 4.9. ∎
@@
For all $\varepsilon>0$, there exists $\tau>0$, so that for all $\eta_{1}>0$, there exists $\eta,\delta_{0}>0$ so that the following holds for all $\delta\in(0,\delta_{0}]$. Let $\beta\geq 2$ and let $(\mathbb{T},Y)_{\delta}$ be a set of $\delta^{-\beta}$ essentially distinct $\delta$-tubes that satisfy the Convex Wolff Axioms with error $\delta^{\eta}$, with $Y$ a $\delta^{\eta}$-uniformly dense shading. Then at least one of the following must be true:
@@
    $\bigcup_{\mathbb{T}}Y(T)$ has discretized Assouad dimension at least $3-\varepsilon$ at scale $\delta$, with scale separation $\delta^{-\eta}$.
@@
    There exists $\rho\in[\delta,\delta^{\varepsilon}]$ and a set $\tilde{\mathbb{T}}$ of $\rho$-tubes, with the following properties:
@@
      $\tilde{\mathbb{T}}$ satisfies the Convex Wolff Axioms with error $\rho^{-\eta_{1}}$.
@@
    There exists $\rho\in[\delta^{1-\varepsilon},1]$, a $\rho$-tube $T_{\rho}$, and a set $\tilde{\mathbb{T}}\subset\mathbb{T}[T_{\rho}]$, so that the following holds:
@@
Decreasing $\eta_{1}$ if necessary, we may may suppose that $\eta_{1}\leq 1$. Fix a number $\eta_{2}\leq\eta_{1}\varepsilon/40$ of the form $\eta_{2}=1/N>0$. Apply Lemma 4.9 with $\eta_{2}$ in place of $\varepsilon$. If Conclusion (A) of Lemma 4.9 holds, then Conclusion (A) of Proposition 4.1 holds (since $\eta_{2}\leq\varepsilon$) and we are done. If instead Conclusion (B) of Lemma 4.9 holds, then there exists:
  • A set $\mathbb{T}^{\prime}\subset\mathbb{T}$,
  • For each number $\rho$ of the form $\delta^{i/N}$, a uniform partitioning cover $\mathbb{T}_{\rho}$ of $\mathbb{T}^{\prime}$.
These objects have the properties described in Conclusion (B) of Lemma 4.9.
Suppose there exists a scale $\rho\in[\delta,\delta^{\varepsilon}]$ of the form $\rho=\delta^{i/N}$ so that #$\#\mathbb{T}_{\rho}\geq\delta^{-\beta-\tau}$. Then by Remark 2.11, $\mathbb{T}_{\rho}$ satisfies the Convex Wolff Axioms with error $\delta^{-2\eta}\leq\rho^{-2\eta/\varepsilon}\leq\rho^{-\eta_{1}}$. Thus Conclusion (C) holds, with $\tilde{\mathbb{T}}=\mathbb{T}_{\rho}$.
Henceforth we shall suppose that for each scale $\rho\in[\delta,\delta^{\varepsilon}]$ of the form $\rho=\delta^{i/N}$, we have
#$\#\mathbb{T}_{\rho}\leq\delta^{-\beta-\tau}.$ (4.37)
Step 2. Our goal in this step is to show that either Conclusion (D) of Proposition 4.1 holds, or for every $\rho\in[\delta,\delta^{\varepsilon}]$ of the form $\rho=\delta^{i/N}$, we have
#$\#\mathbb{T}_{\rho}\geq\rho^{-\beta+\varepsilon/3},\quad\rho\in[\delta,\delta^{\varepsilon}].$ (4.38)
First, we consider the case where $\beta\leq 2+\varepsilon/4$; in this case (4.38) will always hold. Indeed, let $\rho\in[\delta,\delta^{\varepsilon}]$ be of the form $\rho=\delta^{i/N}$. Then for each $T_{\rho}\in\mathbb{T}_{\rho}$, we have
#$\frac{\#\mathbb{T}^{\prime}}{\#\mathbb{T}_{\rho}}=\#\mathbb{T}^{\prime}[T_{\rho}]\leq\delta^{-2\eta}\rho^{2}\#\mathbb{T}^{\prime},$
where the second inequality follows from the fact that $\mathbb{T}^{\prime}$ satisfies the Convex Wolff Axioms with error $\delta^{-2\eta}$. We conclude that #$\#\mathbb{T}_{\rho}\geq\delta^{2\eta}\rho^{-2}$. In particular, if $\beta\leq 2+\varepsilon/4$, and if we select $\eta$ sufficiently small, then
#$\#\mathbb{T}_{\rho}\geq\delta^{2\eta}\rho^{-2}\geq\rho^{2\eta/\varepsilon+\varepsilon/4}\rho^{-\beta}\geq\rho^{-\beta+\varepsilon/3},$ (4.39)
@@
Next we consider the case where $\beta>2+\varepsilon/4$. Suppose there exists some $\rho\in[\delta,\delta^{\varepsilon}]$ of the form $\rho=\delta^{i/N}$ for which #$\#\mathbb{T}_{\rho}<\rho^{-\beta+\varepsilon/4}.$ Let $\rho\in[\delta,\delta^{\varepsilon}]$ be the smallest number of the form $\delta^{iN}$ for which this is true. Note that $\rho\geq\delta^{1-\varepsilon/20}$; this is because the tubes in $\mathbb{T}^{\prime}$ are essentially distinct, and thus #$\#\mathbb{T}_{\rho}\geq(\delta/\rho)^{4}(\#\mathbb{T}^{\prime})\geq(\delta/\rho)^{4}\delta^{-\beta+2\eta}$.
Let $T_{\rho}\in\mathbb{T}_{\rho}$. By Conclusion (B) of Lemma 4.9, for each $s\in[\delta,\rho]$ of the form $\delta^{jN}$, each $t\in[s,\rho]$, and each $s\times t\times 1$ rectangular prism $R\subset T_{\rho}$, we have
#$\#\{T\in\mathbb{T}^{\prime}[T_{\rho}]\colon T\subset R\}$ $\leq\delta^{-\eta_{2}}\frac{t}{s}\frac{(\#\mathbb{T}^{\prime})}{(\#\mathbb{T}_{s})}\leq\delta^{-\eta_{2}}\frac{t}{s}\frac{\delta^{-\beta}}{\rho^{-\beta+\varepsilon/4}}\Big{(}\frac{s}{\rho}\Big{)}^{\beta-\varepsilon/4}$
@@
where the second inequality used the fact that #$\#\mathbb{T}_{s}\geq s^{-\beta+\varepsilon/4}$ (this follows from the minimality of $\rho$); the third inequality used the assumption $\beta-\varepsilon/4\geq 2$; and the fourth inequality used the assumptions #$\#\mathbb{T}^{\prime}\geq\delta^{-\beta+\eta_{1}}$ and #$\#\mathbb{T}_{\rho}\leq\rho^{-\beta+\varepsilon/4}$.
@@
Define $\tilde{\mathbb{T}}=\mathbb{T}^{\prime}[T_{\rho}]$, and let $W\subset T_{\rho}$ be a convex set that contains at least one tube from $\tilde{\mathbb{T}}$. Then $W$ is comparable to a $s\times t\times 1$ prism, for some $\delta\leq s\leq t\leq\rho$; increasing $s$ (and possibly $t$) by a factor of at most $\delta^{-1/N}$, we may suppose that $s$ is of the form $\delta^{jN}$. Thus we have
@@
#$\#\{T\in\tilde{\mathbb{T}}\colon T\subset W\}\lesssim\delta^{-4\eta_{2}}\frac{|W|}{|T_{\rho}|}(\#\tilde{\mathbb{T}})\leq(\delta/\rho)^{-40\eta_{2}/\varepsilon}\frac{|W|}{\rho^{2}}(\#\tilde{\mathbb{T}})\leq(\delta/\rho)^{-\eta_{1}}\frac{|W|}{\rho^{2}}(\#\tilde{\mathbb{T}}).$
@@
Thus our set $\tilde{\mathbb{T}}$ satisfies Conclusion (D) from Proposition 4.1, and we are done. Henceforth we shall assume that (4.38) holds.
@@
Step 3. At this point, we have reduced to the case where (4.37) and (4.38) hold for all $\rho\in[\delta,\delta^{\varepsilon}]$ of the form $\rho=\delta^{i/N}$. We will show that Conclusion (B) holds. To begin, note that Items (i) and (iii) from Definition 2.15 hold for $\mathbb{T}^{\prime}$: for each $\rho_{0}\in[\delta,\delta^{\varepsilon}]$, we can select $\rho\in[\rho_{0},\delta^{-\varepsilon}\rho_{0}]$ of the form $\rho=\delta^{i/N}$; $\mathbb{T}_{\rho}$ is a partitioning cover of $\mathbb{T}^{\prime}$, so Item (i) holds, and Item (iii) holds by (4.38) (if $\rho_{0}>\delta^{\varepsilon}$ then we can select $\rho=1$ and there is nothing to prove).